Geometry. The length of a rectangle is 2 in. more than twice its width. If the
perimeter of the rectangle is 34 in., find the dimensions of the rectangle.

Answer by SHUgrad05(58)

Let w=width and l=length.
Perimeter of a rectangle is P=w+w+l+l
length=2w+2
P=w+w+2w+2+2w+2=34
6w+4=34
6w=30
w=5
l=2(5)+2=12
width=5 in
length=12 in
Check: Perimeter=2w+2l
10+24=34 in